OpenStudy (anonymous):

HELP! MEDAL REWARD!! Consider the line y=-1/4x+6. What is the slope of a line parallel to this line? What is the slope of a line perpendicular to this line?

OpenStudy (anonymous):

slope of parallel lines are same ..

OpenStudy (anonymous):

so what would the answers be? lol

OpenStudy (anonymous):

y=-1/4x+6. slope of the given line ( form y= mx+c) = -1/4 so slope of line parallel to this will also be -1/4 and for the line perpendicular to the given line will be 4 ( neagtive reciprocal of -1/4)